Compound X contains only carbon , hydrogen and oxygen . A 0.2000 g sample of Compound X was burnt completely to yield 0.3906 g of CO2 and 0.200 g of H2O . Determine the empirical formula of Compound X .

1
Expert's answer
2022-04-14T14:37:03-0400

CO2

Molar mass – 44.01 g/mol

44.01 g of CO2 contains 12.01 g of carbon (1 mol).

0.3906 * 12.01 / 44.01 = 0.1066 g C

0.3906 / 44.01 = 0.008875 mol C

 

H2O

Molar mass – 18.02 g/mol

18.02 g of H2O contains 2.02 g H (2 mol)

0.200 * 2.02 / 18.02 = 0.0224 g H

0.200 * 2 / 18.02 = 0.0222 mol H

 

Total mass C and H: 0.1066 g C + 0.0224 H = 0.129 g

Δ = 0.2000 – 0.129 = 0.071 g – mass of O

Moles O = 0.071 / 16.00 = 0.00444 mol

The mole ratio – C : H : O = 0.008875 : 0.0222 : 0.00444 = 2 : 5 : 1

C2H5O